OpenStudy (anonymous):

which of the following are solutions to the equation below? Check all that apply. 4x^2+4x+1=9

OpenStudy (anonymous):

A.\[x=\sqrt{10}-\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

B.x=1

OpenStudy (anonymous):

C.\[x=\sqrt{2}\] d.\[x=-\sqrt{10}-\frac{ 1 }{ 2 }\] E.x=-2 F.x=square root of 2 over 2

OpenStudy (anonymous):

im pretty sure it B and E

OpenStudy (anonymous):

YUP! x = -2 x = 1

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

which of the following are solutions to the equation below? Check all that apply. (x+3)^2=10

OpenStudy (anonymous):

a. x=\[x=10+\sqrt{3}\]

OpenStudy (anonymous):

B.\[x=\sqrt{7}\]

OpenStudy (anonymous):

C.\[x=-\sqrt{7}\]

OpenStudy (anonymous):

D.\[x=\sqrt{10}-3\]

OpenStudy (anonymous):

E.\[x=10-\sqrt{3}\]

OpenStudy (anonymous):

F.\[x=-\sqrt{10}-3\]

OpenStudy (anonymous):

i think its D. and F.

OpenStudy (anonymous):

USE DIFFERENT QUESTIONS \[x = -3-√10\]\[x = √10 -3\]yup

OpenStudy (radar):

I think so too.

OpenStudy (anonymous):

For this equation are the answers D and F (x+3)^2=10

OpenStudy (anonymous):

yes.. but you need to separate your questions